Web Analytics

A bowl contains 25 balls numbered 1 to 25. A ball is drawn and its number is noted. Without replacing the first ball, another ball is drawn.

The probability that the numbers on both balls are odd numbers is

Verified Answer :

To solve the question, first find the number of  odd numbers from 1 to 25 and they are:(1,3,5,7,9,11,13,15,17,19,21,23,25)=13 There are 13 odd numbers from 1 to 25.
The probability that the first number is an odd is 13/25 
Since the first ball wasn't placed back, 24 balls were left. There are 12 odd numbers from 1 -24
so the probability that the second number is odd is 12/24=1/2
13/25+1/2=51/52